OpenStudy (lxelle):

Show that the equation sin(x-60) - cos(30-x) = 1 can be written in the form cosx = k, where k is a constant.

OpenStudy (michele_laino):

Please use these identities: \[\begin{gathered} \sin \left( {x - 60} \right) = \sin x\cos 60 - \cos x\sin 60 = \frac{1}{2}\left( {\sin x - \sqrt 3 \cos x} \right) \hfill \\ \cos \left( {30 - x} \right) = \cos 30\cos x + \sin 30\sin x = \frac{1}{2}\left( {\sqrt 3 \cos x + \sin x} \right) \hfill \\ \end{gathered} \]

OpenStudy (lxelle):

Did alrd, i couldnt continue.

OpenStudy (lxelle):

i got -sqrt3cos

OpenStudy (michele_laino):

that's right! the left side is : \[ - \sqrt 3 \cos x\]

OpenStudy (lxelle):

then?

OpenStudy (lxelle):

it's okay.

OpenStudy (michele_laino):

then you can write: \[\begin{gathered} - \sqrt 3 \cos x = 1 \hfill \\ \cos x = - \frac{1}{{\sqrt 3 }} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok